OpenStudy (anonymous):

Find Exponetial Form. (5)√(7x^(-1))

OpenStudy (anonymous):

\[\frac{5}{\sqrt{7^{-1}}}\] Let's start by rewriting the square root as a fraction. Taking the square root of something is the same as raising it to the 1/2 power. \[\frac{5}{(7^{-1})^\frac{1}{2}}\] We can multiply exponents to get \[\frac{5}{7^{\frac{-1}{2}}}\] Next, we can make a negative exponent into a positive exponent by moving it to the top/bottom. \[5(7^{\frac{1}{2}})\]